Hi
hazel had money to buy 25 chocolate cookies, each choc cookie costs as much as 2 butter cookies.hazel borrowed another 6 dollars to buy 20 choc and 20 butter cookies. How much did hazelinitially have.
Thanks

Found 2 solutions by ikleyn, advanced_Learner:
Answer by ikleyn(52905)

.
Hi
hazel had money to buy 25 chocolate cookies, each choc cookie costs as much as 2 butter cookies.hazel borrowed
another 6 dollars to buy 20 choc and 20 butter cookies. How much did hazel initially have.
Thanks
~~~~~~~~~~~~~~~~~~~~~

Hazel initially had money to buy 25 chocolate cookies, or, equivalently, 50 butter cookies.

After borrowing 6 dollars she buys 20 choc and 20 butter cookies that are equivalent to 2*20 +20 = 60 butter cookies.

Hence, 60-50 = 10 butter cookies cost 6 dollars.

Then 50 cookies cost 5*6 = 30 dollars.

It is the amount Hazel had initially  (see my uppermost line).

Answer.  Hazel did have 30 dollars initially.


Answer by advanced_Learner(501)

25 cholate----X
5 chocolate--6